A 3.5 L sample at 360C and 5.2 atm contains 0.7 moles

   of a gas. If 0.3 mole of the gas is added to the sample

   without changing the temperature and pressure, what is

   the new volume?


1
Expert's answer
2021-04-30T04:51:08-0400

If 0.3 mole is added, then total mole = 0.3+0.7=1

PV=nRT

V=nRT/P

= 1 × 0.082 × 633 / 5.2

=9.98L